OpenStudy (monkey*):

Which expressions are equivalent to 1/9*1/9*1/9 ? (1/9)^-3 9^5/9^2 None of the above

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

correct answer is (1/9)^3

OpenStudy (maddy1251):

\[\frac{ 1 }{ 9 } x \frac{ 1 }{ 9 } x \frac{ 1 }{ 9 } = 0.0013717...\]

OpenStudy (maddy1251):

Ah those were suppose to be multiplying symbols, not variables. Still, same idea as multiplication. :P Now lets see what the answer choices equal.

OpenStudy (maddy1251):

A) = 729, when 1/9^(-3) = 729, same as 9^(3)

OpenStudy (maddy1251):

OS is being slow so sorry for the delay. :(

OpenStudy (maddy1251):

I am assuming C) is 9^(5) / 9^2 = 59049 / 81 = 729 as well.

OpenStudy (maddy1251):

So I agree with oncer, none of the above.

OpenStudy (monkey*):

Thank you!

OpenStudy (luckyness):

Great job @Maddy1251 !! :)

OpenStudy (maddy1251):

@Luckyness Why thank you!! I try to show the work the best I can to explain answers, rather ran DA. So thanks ^.^

OpenStudy (luckyness):

:) I appreciate it!